OpenStudy (anonymous):

Which of the following equations has only one solution? x^2 - 8x + 16 = 0 x(x - 1) = 8 x^2 = 16

OpenStudy (anonymous):

x^2-8x+16=0

OpenStudy (anonymous):

put them in standard form and calculate b^2-4ac if it is zero, then the equation has only one solutions (actually two equal roots)

OpenStudy (anonymous):

the first one quadratic formula

OpenStudy (radar):

\[-(-8)\pm \sqrt{(-8)^{2}-4(16)}\over 2\]\[8\pm \sqrt{0}\over 2\]

OpenStudy (anonymous):

i guess we do not have to actually calculate the roots for this question.... just calculating b^2-4ac will give us the answer....

OpenStudy (anonymous):

You can also factor the first one...you'll get (x-4)(x-4)=0 and x can only equal 4

OpenStudy (radar):

True, but was just checking it out, what does a discriminate = zero signify?

OpenStudy (anonymous):

Only one solution. (Because there's only one square root of zero).

OpenStudy (anonymous):

if D = 0, two equal roots i.e we can say one solution if D>0, two distinct real roots if D<0, imaginary roots

OpenStudy (radar):

Yes, my book says one (repeated) solution, but in this problem, I don't see a repeated solution, do you?

OpenStudy (anonymous):

How come nobody is helping Nicolette?

OpenStudy (anonymous):

just get a ti 84 plus and it will do it for you

OpenStudy (radar):

I guess when it is identical it is difficult to discern lol

OpenStudy (anonymous):

see it says two roots in the sense (8+0)/2 and (8-0)/2 but both result in 4.... so one solution

OpenStudy (anonymous):

Because he/she is posting boring financial questions.

OpenStudy (anonymous):

lol

OpenStudy (radar):

Have a nice day all.

OpenStudy (anonymous):

same to u radar.....

OpenStudy (anonymous):

a

OpenStudy (anonymous):

\[x ^{2}-8x+16\]